OpenStudy (anonymous):

Sally gave her brother $3.95 in quarters and dimes. If the number of quarters is 6 more than the number of dimes, how many quarters did Sally give her brother? Sally gave her brother quarters

OpenStudy (anonymous):

just the answer please or something close

OpenStudy (andrew8270):

well let quarters be x and dimes be y. since there is 6 more quarters... y+6=x... y dimes are worth 0.1dy x quarters are worth 0.25x The problem states: "$3.95 in quarters and dimes" That means 0.1d + 0.25q = 3.95 That is our second equation. Here are the two equations: q = d + 6 0.1d + 0.25q = 3.95

OpenStudy (andrew8270):

sorry I changed x to q and y to d to make more sense